Consider the titration of 50.0 mL of 0.0200 M In+ by 0.0500 m Cu2+ using Pt and

saturated calomel electrodes to find the end point.

a. Write a balanced titration reaction.

b. Write two different half-reactions for the indicator electrode.

c. Write two different Nernst equations for the net cell reaction.

d. Calculate E at the following volumes of Cu2+: 1.00, 20.0, 39.0, 40.0, 41.0, and 80.0

mL. Sketch the titration curve.

2. Consider the titration of 50.0 mL of 0.0250 M Fe3+ by 0.0500 M Cu+ to give Fe2+ and Cu2+ ,

using Pt and saturated AgAgCl electrodes to find the end point.

a. Write a balanced titration reaction.

b. Write two different half-reactions for the indicator electrode.

c. Write two different Nernst equations for the net cell reaction.

d. Calculate E at the following volumes of Cu2+: 1.0, 12.5, 24.5, 25.0, 25.5 and 30.0

mL. Sketch the titration curve 


1
Expert's answer
2021-05-31T02:24:52-0400

1)


a)  Balanced titration reaction


"In^{+} + Cu^{2+}\u21cb In^{2+}+Cu^+"


b) half-reaction


"In^{+} \rightarrow In^{2+}+ e^-" Eo = 0.4 V


"Cu^{2+} + e^- \rightarrow Cu^+" Eo = 0.161 V


c)Nernst equations for half reactions



"E=0.4-0.059log({[In^{2+ }]\over[In^{+}]})"


"E=0.161-0.059log({[Cu^{+ }]\over[Cu^{2+}]})"



2)


a)  Balanced titration reaction


"Fe^{3+} + Cu^+ \rightarrow Fe^{2+} + Cu^{2+}"


b)half-reaction


"Cu^{2+}+ e^- \rightleftharpoons Cu^+" Eo = 0.161 V


"Fe^{3+}+ e^- \rightleftharpoons Fe^{2+}" Eo = 0.767 V


c) Nernst equations for half reactions



"E=0.767-0.059log({[Fe^{2+ }]\over[Fe^{3+}]})"


"E=0.161-0.059log({[Cu^{+ }]\over[Cu^{2+}]})"